iuytyuioiuytyuiop:

The length of a rectangular garden is 4 feet longer than the width. If the perimeter is 192 feet, what is the area of the garden? Do not include units in your answer.

iuytyuioiuytyuiop:

@AZ

AZ:

So first the formula Perimeter of a rectangle = 2 * length * width Now your question says length is 4 feet longer than the width can you write that sentence as an equation?

iuytyuioiuytyuiop:

2 * 4 *192

AZ:

the perimeter is 192 but first we need to find the length and width all we know is the relationship between length and width the length equals to 4 ft more than width how would we translate this into some math with letter like L and W

iuytyuioiuytyuiop:

is the length 4?

iuytyuioiuytyuiop:

4 * w

iuytyuioiuytyuiop:

Okay I see now

AZ:

4 feet MORE NOT 4 times the width so it's just addition L = 4 + w

AZ:

so remember the first formula I mentioned? Perimeter = 2 * (length + width) now we have to plug in length = 4 + width and we'll get perimeter = 2*(4 + width + width) and let's just use the variable w for width perimeter = 2*(4 + w + w) remember perimeter = 192 so can you solve for w?

iuytyuioiuytyuiop:

192 divided by 2?

AZ:

so if we do that, we'll get 192/2 = 4 + w + w what is 192/2 =

iuytyuioiuytyuiop:

96

AZ:

so we have 96 = 4 + w + w what is w + w = ?

iuytyuioiuytyuiop:

I don't know what number w is or how to find it

AZ:

it's just a variable so if I have one apple + one apple then I have 2 apples just like that x + x = 2x so what is w + w =

iuytyuioiuytyuiop:

2w

AZ:

exactly! so we have 96 = 4 + 2w we want to take the 4 to the other side, so what do we do? Do we have to add 4 on both sides or subtract 4 on both sides?

iuytyuioiuytyuiop:

subtract 4 on both sides

AZ:

good! so what do we get?

iuytyuioiuytyuiop:

do we subtract 4 from 96 and 2w

AZ:

we subtract 4 on both sides so like 96 - 4 = 4 + 2w - 4 and so on the right side, the 4 and -4 would cancel out giving us 0 so we're left with 96 - 4 = 2w does that make sense?

iuytyuioiuytyuiop:

yep

iuytyuioiuytyuiop:

do we keep subtracting

AZ:

we don't keep subtracting haha so what is 96 - 4 =

iuytyuioiuytyuiop:

92

AZ:

good so now we don't have to subtract anymore 92 = 2w so to find w, divide by 2 on both sides so that we can get w all by itself

iuytyuioiuytyuiop:

46 = 1w

iuytyuioiuytyuiop:

46=w

AZ:

Good! So now that we know w, we can find the length remember that sentence we wrote as a math expression/equation where length was 4 feet more than the width so length = 4 + width what is the length if our width is 46

iuytyuioiuytyuiop:

11.5

AZ:

mhmmm not sure how you got that length = 4 + width width is 46 length = 4 + 46 length = ??

iuytyuioiuytyuiop:

50

AZ:

yes so now we know length is 50 feet width is 46 feet we want to find the area of the rectangle area = length * width can you find the area

iuytyuioiuytyuiop:

2,300

snowflake0531:

yepp

AZ:

Yes!